Syllabus UTK – M351 – Honors Calculus II Spring 2004, Jochen Denzler, MTWF 1:25–2:15, Ayres 309A

Textbook: any halfway intelligent Calculus textbook (none is really excessively intelligent); old editions fine.

I have some Calc’2 material posted on my website, from 3 years ago at Notre Dame. Ignoring course-specific comments therein, I’ll basically follow them in style and contents, to the extent that they cover the material.

Course contents: Integrals and applications (volume, arclength, maybe surface area of rotation bodies); practical evaluation of integrals; exponential, logarithm, hyperbolic functions; sequences and series, in particular Taylor series.

Homework: Will usually not consist of mere routine or drill problems, but require some effort and thought. It is perfectly ok to come in with questions. Approximately one class per week (probably Thursday, but we’ll be flexible) will be mainly discussion and presentation of homework.

Office hours: In order to cope with a weakness in time management, I intend to reserve Thursdays for research. Otherwise, I do accomodate drop-in whenever possible and you may also schedule appointments.

Regular office hours will be posted later. My office is Ayres 317 E, phone 4-5325. Email is denzler@math.utk.edu, but I may not read it for half a day or for an entire weekend. My website is http://www.math.utk.edu/˜ denzler.

Grade: Today, I suggest the following: We’ll have three in-class exams and a comprehensive final. The final counts 35%, homework 20%, the in-class exams each 15%. In the small size honors class we are, this rule is up to negotiation, and if we agree otherwise, you’ll get an amended syllabus by the end of the week.